XYZ Corporation has a deferred compensation plan under which it allows certain employees to defer up to 40 percent of their salary for five years. For purposes of this problem, ignore payroll taxes in your computations.

b. Assume Julie, an XYZ employee, has the option of participating in XYZ’s deferred compensation plan. Julie’s marginal tax rate is 37 percent and she expects the rate to remain constant over the next five years. Julie is trying to decide how much deferred compensation she will need to receive from XYZ in five years to make her indifferent between receiving the current salary of $13,200 and receiving the deferred compensation payment. If Julie takes the salary, she will invest it in a taxable corporate bond paying interest at 7 percent annually (after taxes). What amount of deferred compensation would accomplish this objective?
Answer #1

ANSWER

The amount Julie would receive if she decides to take the salary now is calculated as follows:

Amount Received Now = Current Salary*(1-Tax Rate) = 13,200 *(1-40%) = $7,920

This amount of $7,920 will be invested by Julie in taxable corporate bonds. The value after 5 years is calculated as follows:

Value after 5 Years = Amount*(1+Interest Rate)^Years = 7,920*(1+7%)^5 = $11,108.21

____

The amount of deferred compensation that will make the Julie indifferent is determined below:

Amount of Deferred Compensation = Value after 5 Years/(1-Tax Rate) = 11,108.21/(1-37%) = $17,632.07 or $17,632(answer)

_____________________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU**************